OpenStudy (anonymous):

Find the derivative of f(x) =(-9/x) at x = -8. i dont know how to solve it as a fraction?

OpenStudy (cggurumanjunath):

f'(x)=9/(x^2) f'(-8)=9/((-8)^2)=9/64

OpenStudy (anonymous):

it's the same as (-9) * (1/x)

OpenStudy (anonymous):

thanks!(:

OpenStudy (cggurumanjunath):

you are welcome.

OpenStudy (cggurumanjunath):

ask me more questions on differentiation.